OpenStudy (anonymous):

Write the equation of the line that passes through (–3, 4) with a slope of 2 in point-slope form.

OpenStudy (anonymous):

y-4=2(x+3)

OpenStudy (anonymous):

point slope form is: y-y1=m(x-x1) you plug in the slope where the m is, and the point #s in the x1 and y1 spots.

OpenStudy (anonymous):

equation of required line is y-y1=m(x-x1) where (x1,y1) is your point through which line is passing, and 'm' is slope of line y-3=2(x-(-3)) y-3=2x + 6 2x-y +9=0

OpenStudy (anonymous):

@prashant- you put in a 3 instead of a 4, and i think he isn't supposed to simplify, just keep it in point slope form

OpenStudy (anonymous):

its an honest mistake just make it as y-4=.......then everything is same... thanks sunyah

OpenStudy (anonymous):

yup yup of course! :)